OpenStudy (anonymous):

lim (x^2-1)/(x-1) x->1

OpenStudy (chaise):

I don't think it is infinite.

OpenStudy (anonymous):

answer in the back of the book is 2 can you explain

OpenStudy (chaise):

(x^2-1) = (x+1)(x-1) (x+1)(x-1)/(x-1) The x-1's cancel off. Your left with x+1 Sub in your value. 2 is the answer

OpenStudy (chaise):

The limit is 2.

OpenStudy (anonymous):

limit is 2...

OpenStudy (anonymous):

The answer is 2. You factor the numerator so it is (x+1)(x-1)... (x-1) cancels with the (x-1) in the denominator. As x approaches closer to 1, the y will be 2. Therefore, the limit is 2.

OpenStudy (anonymous):

Okay I didn't factor the numerater and was getting 0

OpenStudy (anonymous):

if you substitute and the answer results to 0/0, you can factor out..

OpenStudy (anonymous):

do you always factor if there is a exponent

OpenStudy (anonymous):

You factor if you can... lol

OpenStudy (anonymous):

you can also derivate (x^2 - 1) and (x - 1) getting lim 2x/1, x -> 1

OpenStudy (dumbcow):

you can try L'hopitals rule will work here as well \[\frac{(d/dx) x^{2}-1}{(d/dx) x-1} = 2x\]

OpenStudy (anonymous):

Thanks guys appreciate it